A force of 50.7 Newtons pulls due west on a structure, and a second force of 81.3 Newtons pulls S 21o W. Find the principle angle of the resultant.

Round the answer to 2 decimal places.
1
Expert's answer
2018-09-19T12:18:08-0400
R_x=v+u_x=50.7+81.3 cos⁡21=126.6 N.
R_y=u_y=81.3 sin⁡21=29.14 N.
The principle angle of the resultant is
θ=tan^(-1)⁡〖R_y/R_x 〗=tan^(-1)⁡〖(29.14 )/126.6〗=S 12.96° W